# supReferences

### v4

@@ A.2.1 A simple but useful technical lemma
Let $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$. If $r=0$ the result of the lemma is trivial and we therefore assume that $r\neq 0$ in what follows. Then,
@@
showing the first part of the lemma. On the other hand, using the mean value theorem and noting that $|\Phi^{\prime}(x)|=|\varphi(x)|\leq 1/\sqrt{2\pi}$ for all $x\in\mathbb{R}$, we have
@@
A.2.2 A first useful property of the model of Section 2.1
@@
$c_{k}=\frac{(c+c_{k-1})\rho^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\rho_{k}=\prod_{i=0}^{k-1}\frac{\rho}{1+(c+c_{i})\sigma^{2}},\quad\sigma_{k}^{2}=\sigma_{k-1}^{2}+\frac{\rho_{k-1}^{2}\sigma^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\forall k\geq 1.$
Then, for all $k\geq 1$, we have
@@
The result in the first part of the lemma can be deduced from the calculations in Del Moral and Jacod, (2001) while the convergence result for the sequence $(\sigma_{k})_{k\geq 1}$ can be directly obtained from Del Moral and Horton, (2023, Theorem 3.5). However, for sake of completeness the whole lemma is proved in what follows.
@@
so that $c_{k}=f(c_{k-1})$ for all $k\geq 1$. Easy computations show that the function $f$ is non-decreasing on $[0,\infty)$, and thus the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and such that $c\rho^{2}/(1+c\sigma^{2})\leq c_{k}\leq\rho^{2}/\sigma^{2}$ for all $k\geq 1$.
@@
To proceed further we remark that, since the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and bounded, the sequence converges in $[0,\infty)$ and we let $c_{\infty}=\lim_{k\rightarrow\infty}c_{k}$. In addition, simple calculations show that
@@
is the unique value of $x\in[0,\infty)$ such that $x=f(x)$, and since we are assuming that $|\rho|\geq 1+c\sigma^{2}>1$ the constant $c_{\star}$ is such that
$c_{\star}\geq\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}}}{2\sigma^{2}}=\frac{\rho^{2}-(1+c\sigma^{2})}{\sigma^{2}}>\frac{\rho-(1+c\sigma^{2})}{\sigma^{2}}.$ (16)
On the other hand, since $f$ is continuous on $[0,\infty)$ we have
@@
showing that $c_{\infty}=c_{\star}$, and thus $\lim_{k\rightarrow\infty}c_{k}=c_{\star}$. Together with (16), this implies that there exists a $k_{\star}\in\mathbb{N}$ such that
$c_{k_{\star}}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}$
and thus, using (15) and the fact that the sequence $(c_{k})_{k\geq 0}$ is non-decreasing, we have
@@
$|\sigma_{m}^{2}-\sigma_{k}^{2}|=\Big{|}\sum_{s=k}^{m-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}\Big{|}\leq\frac{\sigma^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\rho_{s}^{2}\leq\frac{\sigma^{2}C^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\epsilon_{\star}^{2s}\leq\epsilon_{\star}^{2k}\frac{\sigma^{2}C^{2}}{(1-\epsilon_{\star})(1+c\sigma^{2})}.$ (17)
Using (17) it is easily verified that the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is Cauchy and therefore converges in $[0,\infty)$, and we let $\sigma^{2}_{\infty}=\lim_{k\rightarrow\infty}\sum_{s=0}^{k-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}$. Remark that since the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is non-decreasing and such that $\sigma^{2}_{k}\geq\sigma^{2}/(1+c\sigma^{2})$ it follows that $\sigma^{2}_{\infty}>0$, and remark that, by (17), we have

### v5

@@ 5.2.1 A simple but useful technical lemma
Let $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$. If $r=0$ the result of the lemma is trivial and we therefore assume that $r\neq 0$ in what follows. In this scenario, we have
@@
showing the first part of the lemma. On the other hand, using the mean value theorem and noting that $|\Phi^{\prime}(x)|=|\varphi(x)|\leq 1/\sqrt{2\pi}$ for all $x\in\mathbb{R}$, we obtain that
@@
5.2.2 A general 1-dimensional Koksma-Hlawka inequality
The next result is a general $s=1$ dimensional Koksma-Hlawka inequality, where the integration is w.r.t. an arbitrary distribution $\mu\in\mathcal{P}(\mathbb{R})$. By contrast, in the original Koksma-Hlawka inequality, the dimension $s\geq 1$ is arbitrary but the integration is w.r.t. the $\mathcal{U}(0,1)^{s}$ distribution (see e.g. Dick and Pillichshammer,, 2010, Proposition 2.18, page 33). It is also worth mentioning that Aistleitner and Dick, (2015) derived a version of the Koksma-Hlawka inequality where the integration is w.r.t. an arbitrary probability distribution on $(0,1)^{s}$ for some arbitrary $s\geq 1$.
Lemma 4.
Let $\infty\leq a<b\leq\infty$ and let $f:\mathbb{R}\rightarrow\mathbb{R}$ be continuously differentiable on the interval $(a,b)$ and such that $V(f\mathbbm{1}_{(a,b)})<\infty$, and let $\pi,\mu\in\mathcal{P}(\mathbb{R})$. Then, with the convention that $f(a)=0$ when $a=$-$\infty$ and with the convention that $f(b)=0$ when $b=\infty$, we have
Proof.
Assume first $(a,b)\in\mathbb{R}^{2}$ and note that, since the function $f$ is continuous, we have $\int_{a}^{b}f(x)\mathrm{d}\big(F_{\pi}(x)-F_{\mu}(x)\big)=\int_{a}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))$. Then, using the integration by part formula for Riemann-Stieltjes integrals (see Apostol,, 1973, Theorem 7.6, page 144), we have
$\begin{split}\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)f^{\prime}(x)\mathrm{d}x&=\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)\mathrm{d}f(x)\\ &=f(b)\big(F_{\pi}(b)-F_{\mu}(b)\big)-f(a)\big(F_{\pi}(a)-F_{\mu}(a)\big)-\int_{a}^{b}f(x)\mathrm{d}\big(F_{\pi}(x)-F_{\mu}(x)\big)\end{split}$
from which we deduce that
$\begin{split}\Big|\int_{a}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))\Big|&\leq|f(b)|\,\big|F_{\pi}(b)-F_{\mu}(b)\big|+|f(a)|\,\big|F_{\pi}(a)-F_{\mu}(a)\big|\\ &+\Big|\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)f^{\prime}(x)\mathrm{d}x\Big|.\end{split}$ (29)
The result of the lemma when $(a,b)\in\mathbb{R}^{2}$ directly follows from (29).
Assume now that $a=$-$\infty$ and that $b\in\mathbb{R}$. Note that under the assumptions of the lemma we have $\sup_{x\in\mathbb{R}}|f(x)|<\infty$ while, by using the reverse Fatou’s lemma, we obtain that $\lim_{x\rightarrow-\infty}F_{\pi}(x)=\lim_{a\rightarrow-\infty}F_{\mu}(x)=0$. Therefore, using (29),
$\Big|\int_{-\infty}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))\Big|$ $=\Big|\lim_{a\rightarrow-\infty}\int_{a}^{b}f(x)\mathrm{d}\big(F_{\pi}(x)-F_{\mu}(x)\big)\Big|$
$\leq\big|f(b)|\,\big|F_{\pi}(b)-F_{\mu}(b)|$
$+\limsup_{a\rightarrow-\infty}\Big|\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)f^{\prime}(x)\mathrm{d}x\Big|$
$\leq\|\pi-\mu\|\Big(|f(b)|+V(\mathbbm{1}_{(-\infty,b)}f)\Big)$
proving the result of the lemma when $a=$-$\infty$ and $b\in\mathbb{R}$.
Using Fatou’s lemma we have $\lim_{x\rightarrow\infty}F_{\pi}(x)=\lim_{a\rightarrow\infty}F_{\mu}(x)=1$, and the result of the lemma for the two remaining case (i.e. for $b=\infty$ and $a\in\mathbb{R}\cup\{-\infty\}$) can proved using similar calculations as above, and is therefore omitted to save space. ∎
5.2.3 A useful property of the model of Section 2.1
@@ 5.2.2 A general 1-dimensional Koksma-Hlawka inequality
$c_{k}=\frac{(c+c_{k-1})\rho^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\rho_{k}=\prod_{i=0}^{k-1}\frac{\rho}{1+(c+c_{i})\sigma^{2}},\quad\sigma_{k}^{2}=\sigma_{k-1}^{2}+\frac{\rho_{k-1}^{2}\sigma^{2}}{1+(c+c_{k-1})\sigma^{2}}$
for all $k\geq 1$. Then, for all $k\geq 1$, we have
@@
The result stated in (30) can be deduced from the calculations in Del Moral and Jacod, (2001) while the convergence result for the sequence $(\sigma_{k})_{k\geq 1}$ can be directly obtained from Del Moral and Horton, (2023, Theorem 3.5). However, for sake of completeness, a proof of the whole lemma is presented below.
@@
so that $c_{k}=f(c_{k-1})$ for all $k\geq 1$. Easy computations show that the function $f$ is non-decreasing on $0,\infty)$, and thus the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and such that $c\rho^{2}/(1+c\sigma^{2})\leq c_{k}\leq\rho^{2}/\sigma^{2}$ for all $k\geq 1$.
@@
To proceed further we remark that, since the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and bounded, the sequence converges in $0,\infty)$ and we let $c_{\infty}=\lim_{k\rightarrow\infty}c_{k}$. In addition, simple calculations show that
@@
is the unique value of $x\in[0,\infty)$ such that $x=f(x)$. Moreover, since we are assuming that $|\rho|\geq 1+c\sigma^{2}>1$, it follows that $\rho^{2}-(1+c\sigma^{2})\geq 0$ and thus the constant $c_{\star}$ is such that
$c_{\star}\geq\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}}}{2\sigma^{2}}=\frac{\rho^{2}-(1+c\sigma^{2})}{\sigma^{2}}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}.$ (34)
On the other hand, since the function $f$ is continuous on $0,\infty)$ we have
@@
showing that $c_{\infty}=c_{\star}$, and thus $\lim_{k\rightarrow\infty}c_{k}=c_{\star}$. Together with (34), this implies that there exists a $k_{\star}\in\mathbb{N}$ such that $c_{k_{\star}}>(|\rho|-(1+c\sigma^{2}))/\sigma^{2}$, and thus, using (33) and the fact that the sequence $(c_{k})_{k\geq 0}$ is non-decreasing, we have
@@
$\begin{split}|\sigma_{m}^{2}-\sigma_{k}^{2}|=\Big|\sum_{s=k}^{m-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}\Big|\leq\frac{\sigma^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\rho_{s}^{2}&\leq\frac{\sigma^{2}C^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\epsilon_{\star}^{2s}\leq\epsilon_{\star}^{2k}\frac{\sigma^{2}C^{2}}{(1-\epsilon_{\star})(1+c\sigma^{2})}.\end{split}$ (35)
Using (35) it is easily verified that the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is Cauchy and therefore converges in $0,\infty)$, and we let $\sigma^{2}_{\infty}=\lim_{k\rightarrow\infty}\sum_{s=0}^{k-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}$. Remark that since the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is non-decreasing and such that $\sigma^{2}_{k}\geq\sigma^{2}/(1+c\sigma^{2})$ it follows that $\sigma^{2}_{\infty}>0$, and remark that, by (35), we have